OpenStudy (anonymous):

A 3.77 g sample of solid KCl (s) is dissolved in 288 mL of water in a coffee cup calorimeter. Once all of the KCl (s) is dissolved in the water, the final temperature of the solution is found to be 12.92°C. If the initial temperature of the water in the calorimeter was 22.12 °C, calculate the calorimeter constant (in J/K) for the coffee cup calorimeter. Report your answer to three significant figures. The heat of solvation of KCl (s) is 17.22 kJ/mol.

OpenStudy (aaronq):

q=m*Cp*dT find q by using the mass of KCl and the heat produced upon solvation.

OpenStudy (anonymous):

i'm getting a negative? i have (17.22)/(3.77)x(12.92-22.12) = -2.076 that doesnt seem right to me

OpenStudy (aaronq):

the heat of solvation should be a negative value because it's an exothermic process

OpenStudy (aaronq):

and you're not using the mass of KCl in the question itself, you're using the mass of the water because thats whats changing temperature.

OpenStudy (anonymous):

right i think i'm a bit confused. So i find the q for water.... (0.288g)(4.18)(12.92-22.12) = -11.075 this is the qsurr? so q sys would be 11.075. but now what do i do?

OpenStudy (aaronq):

you're taking Cp to be the heat capacity for the calorimeter, don't assume Cp for water is 4.183. so find the heat (q) produced from the solvation q=(288 g)*Cp*(-9.2)

OpenStudy (aaronq):

at least thats what i'm getting from it, if you need to take the water and the calorimeter separately then q=(288 g)*(Cp+4.18)*(-9.2)

OpenStudy (anonymous):

right but i dont have the initial temperature for KCl?

OpenStudy (aaronq):

i don't see why you would need it, q = - 17.22 kJ/mol * moles of KCl

OpenStudy (anonymous):

yah i'm lost, haha thanks for trying to help though

OpenStudy (aaronq):

what don't you get?

OpenStudy (anonymous):

what to do with the q of KCl

OpenStudy (aaronq):

you equate it to the calorimeter equation. you're basically first finding the amount of heat produced by the solvation of the ions, you're given the amount in kJ/mol. so convert the mass of KCl to moles then multiply it by the heat of solvation. then use q=288g*Cp*dT